The balance sheet and income statement shown below are for Sneaker Inc. Note that the firm has no amortization charges, it does not lease any assets, none of its debt must be retired during the next 5 years, and the notes payable will be rolled over.

Balance Sheet (Millions of $)
Assets 2019
Cash and securities $ 2,500
Accounts receivable 11,500
Inventories 16,000
Total current assets $30,000
Net plant and equipment $20,000
Total assets $50,000
Liabilities and Equity
Accounts payable $ 9,500
Accruals 5,500
Notes payable 7,000

Total current liabilities

$22,000
Long-term bonds $15,000
Total liabilities $37,000
Common stock $ 2,000
Retained earnings 11,000
Total common equity $13,000
Total liabilities and equity $50,000
Income Statement (Millions of $) 2019
Net sales $87,500
Operating costs except depreciation 81,813
Depreciation 1,531
Earnings bef interest and taxes (EBIT) $ 4,156
Less interest 1,375
Earnings before taxes (EBT) $ 2,781
Taxes 973
Net income $ 1,808
Other data:
Shares outstanding (millions) 500.00
Common dividends $632.73
Int rate on notes payable & L-T bonds 6.25%
Federal plus state income tax rate 35%

Year-end stock price

$43.39

What is the firm's ROE?

What is the firm's profit margin?

What is the firm's P/E ratio?

What is the firm's operating margin?

What is the firm's market-to-book ratio?

What is the firm's inventory turnover ratio?
Answer #1

What is the firm's ROE

=net income/equity

=1808/13000

=13.91%

What is the firm's profit margin

=net income/sales

=1808/87500

=2.07%

What is the firm's P/E ratio

=43.39/(1808/500)

=12.00

What is the firm's operating margin

=EBIT/sales

=4156/87500

=4.75%

What is the firm's market-to-book ratio

=43.39/(13000/500)

=1.67

What is the firm's inventory turnover ratio (using Operating costs except depreciation)

=81813/16000

=5.11

What is the firm's inventory turnover ratio (using sales)

=87500/16000

=5.47